# Hodge structures on conformal blocks

### v1

@@ 5.3. Decomposition complex, following Bakalov-Kirillov
Let $\mathcal{V}$ be a modular functor with associated ribbon category $\mathrm{C}_{\mathcal{V}}$. Then for colors $\lambda_{1},\dotsc,\lambda_{n}\in\Lambda$, we have natural111111with respect to $\mathcal{V}$. isomorphisms:
@@
Let $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ be $2$ modular functors in $\mathfrak{tmod}_{c}$ with respective weakly ribbon categories $\mathrm{C}_{\mathcal{V}_{1}}$ and $\mathrm{C}_{\mathcal{V}_{2}}$. Let $G:\mathrm{C}_{\mathcal{V}_{1}}\rightarrow\mathrm{C}_{\mathcal{V}_{2}}$ be a morphism. For a colored surface $\Sigma=(S,\underline{\lambda})$, and a decomposition $D=(\Gamma,(\phi_{S_{i}})_{i})$ of $S$, we define an isomorphism of $R$-modules $G_{D}:\mathcal{V}_{1}(\Sigma)\rightarrow\mathcal{V}_{2}(\Sigma)$ as follows. For $k=1,2$, the gluing axiom along $\Gamma$ expresses $\mathcal{V}_{1}(\Sigma)$ as a direct sum of tensor powers of terms of the form $\mathcal{V}_{k}(S_{i},\underline{\mu})$ or $\mathcal{V}_{k}(S_{0}^{2},\nu,\nu^{\dagger})$. Now we have isomorphisms:
@@
To prove fullness of $\mathfrak{tmod}_{c}\rightarrow\mathfrak{wRib}$, we will need that $G_{D}$ is independent of the decomposition $D$ (see Section 5.4). For this, by Theorem 5.10, it is sufficient to prove that $G_{D}$ is invariant under $Z$, $B$, $F$ and $S$-moves. So we need to show that if $D$ and $D^{\prime}$ differ by a move, then the associated decompositions of $\mathcal{V}(\Sigma)$ as in Definition 5.11 are related by isomorphisms depending only on $\mathrm{C}_{\mathcal{V}}$ and natural in $\mathcal{V}$. Below, we explain the invariance under $Z$, $B$, $F$-moves. Invariance under $S$-moves will be the subject of the next subsection 5.4.
@@ 5.4. Higher genus and curve operators, following Andersen-Ueno
Andersen and Ueno prove this for $R=\mathbb{C}$. However their proof works without changes for $R$ any field. It consists in showing that a matrix $C$ depending on the data of the modular functor is invertible [2, 5.2]. Assume by contradiction that the statement is false for some $\mathcal{V}$ defined over some ring $R$. Then $\det(C)$ is included in a maximal ideal $\mathfrak{m}$ of $R$. Hence the reduction of $\mathcal{V}$ over $R/\mathfrak{m}$ has non-invertible $C$-matrix. However $R/\mathfrak{m}$ is a field, so this contradicts Andersen and Ueno’s result. ∎
We can now conclude on the fullness of $\mathfrak{mod}_{c}\rightarrow\mathfrak{wRib}$.
@@
Let $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ be two modular functors with respective weakly ribbon categories $\mathrm{C}_{\mathcal{V}_{1}}$ and $\mathrm{C}_{\mathcal{V}_{2}}$. Choose a morphisms $G:\mathrm{C}_{\mathcal{V}_{1}}\rightarrow\mathrm{C}_{\mathcal{V}_{2}}$. Then for each decompositions $D$ of a surface $\Sigma$, Definition 5.11 provides an isomorphism $G_{D}:\mathcal{V}_{1}(\Sigma)\rightarrow\mathcal{V}_{2}(\Sigma)$. If we show that $G_{D}$ is independent of the decompositions $D$, then setting, for each colored surface $\Sigma$, $G_{\Sigma}:=G_{D}$, the family $(G_{\Sigma})_{\Sigma}$ provides an isomorphism of modular functor $\widetilde{G}:\mathcal{V}_{1}\rightarrow\mathcal{V}_{2}$ lifting $G:\mathrm{C}_{\mathcal{V}_{1}}\rightarrow\mathrm{C}_{\mathcal{V}_{2}}$.
@@ 6.1. Higgs bundles and local systems
Let $X$ be a smooth Deligne-Mumford stack. Then we have equivalences between the $\mathbb{C}$-linear symmetric tensor categories 121212i.e. with a tensor product and duals. of:
@@ A Higgs bundle $(\mathcal{E},\theta)$ on a proper smooth DM stack $X$ is of harmonic type if:
Let $\Lambda$ be a set of colors and $r,s\geq 1$ integers. Then the definition of a modular functor of Higgs bundles with level $(r,s)$ is obtained from that of a (non-anomalous) complex modular functor (2.29, 2.42) by replacing the $\mathbb{C}$-local systems $\mathcal{V}_{g}(\underline{\lambda})$ by Higgs bundles $(\mathcal{V}_{g}(\underline{\lambda}),\theta_{g}(\underline{\lambda}))$.
Similarly for genus 00 modular functors and rooted functors.
From the correspondence and its functoriality we immediately get the following.
@@
complex modular functors with semisimple quantum representations (see 3.8);
@@
Similarly for genus 00 modular functors and rooted functors.
@@ 6.2. Systems of Hodge bundles
Hence we can split $\nabla$ into $\overline{\theta}+\partial_{E}+\overline{\partial}_{E}+\theta$ according to the decomposition above. Then $(E,\overline{\partial}_{E},\theta)$ is the associated system of Hodge bundles, and any polarization for $(E,\nabla)$ is also one for $(E,\overline{\partial}_{E},\theta)$. To see this, notice that for any polarization $h$, the metric $K=\pm\sum_{p+q=m}(-1)^{p}h_{\mid E^{p,q}}$ is pluri-harmonic.
Let $(\mathcal{E},\theta)$ be a system of Hodge bundles on $(\mathcal{E},\theta)$. Denote by $E$ the underlying $C^{\infty}$ bundle and by ¯$\overline{\partial}_{E}$ the Dolbeault operator. Choose a polarization $h$. Now let $\partial_{E}$ be the unique $(1,0)$-connection such that $\partial_{E}+\overline{\partial}_{E}$ preserves $h$, and let $\overline{\theta}$ be the adjoint of $\theta$ for $h$. Then $(E,\nabla=\overline{\theta}+\partial_{E}+\overline{\partial}_{E}+\theta)$ is the associated CVHS (with real bidegrees) and $h$ polarizes it. Moreover $\nabla$ is independent of the choice of the polarization $h$. Again, to see this, notice that the metric $K=\pm\sum_{p+q=m}(-1)^{p}h_{\mid E^{p,q}}$ is Hermitian-Yang-Mills.
@@
Let $\mathcal{V}$ be a complex modular functor with semisimple quantum representations. Let $\Lambda$ be its set of colors and $(\mathcal{V},\theta)$ the associated modular functor of Higgs bundles (6.15). Choose a weight function $w$ on $\mathcal{V}$ and some $t\in\mathbb{R}_{+}^{*}\setminus\{1\}$. Then the maps $i$ and $r$ of Lemma 6.20 induce maps:
@@ 7. Ocneanu rigidity
7.1. Deformations of associators and endofunctors of ribbon categories
@@
A ribbon category is a weakly ribbon category (5.5) that is rigid as a monoidal category. In particular, it is fusion category. Ribbon categories over $R$ form a full subcategory $\mathfrak{Rib}$ of $\mathfrak{wRib}$.
Note that our definition of morphisms of fusion and ribbon categories is more restrictive than the usual one.
@@
Similarly, if $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ is a ribbon category over $\mathbb{C}$, a deformation of $c$ is an associator $\widetilde{c}$ such that $(\mathrm{Vect}_{\Lambda}^{{\mathbb{C}[\epsilon]}},\otimes_{{\mathbb{C}[\epsilon]}},\widetilde{c},\beta_{\mathbb{C}[\epsilon]},\theta_{\mathbb{C}[\epsilon]})$ is weakly ribbon. Such a deformation is trivial if there exists a pair $(\mathrm{id}_{\Lambda},(\mathrm{id}+\epsilon\phi))$ as above that is a ribbon isomorphism.
Note that any deformation $\widetilde{c}$ of $c$ is always of the form $c(\mathrm{id}+\epsilon d)$ for $d:(-\otimes-)\otimes-\rightarrow-\otimes(-\otimes-)$ a natural transformation.
@@
We now briefly explain how this result for fusion categories can be extended to ribbon categories. This is folklore, but we take this opportunity to better understand the meaning of Theorem 7.4. Fix a ribbon category $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$. Let $C^{3}(\mathcal{C})$ be the vector space of all natural transformations $d:(-\otimes-)\otimes-\rightarrow-\otimes(-\otimes-)$, let $Z^{3}(\mathcal{C})$ be the subspace of the $d$ such that $c(\mathrm{id}+d)$ is an associator for $(\mathrm{Vect}_{\Lambda}^{{\mathbb{C}[\epsilon]}},\otimes_{{\mathbb{C}[\epsilon]}})$, and let $Z^{3}(\mathcal{C},\beta)$ be the subspace of the $d$ such that $c(\mathrm{id}+d)$ is an associator compatible with the braiding $\beta_{\mathbb{C}[\epsilon]}$.
@@
Ocneanu rigidity (7.4) is equivalent to the statement that $\partial$ is surjective. The map $\partial$ is a differential in the proper Davydov-Yetter complex of $\mathrm{id}_{\mathcal{C}}$. This complex is quasi-isomorphic to the usual Davydov-Yetter complex. Ocneanu rigidity is usually stated as the acyclicity of this latter complex (see for example [13, 7.22, 9.1.3]).
@@
Corollary 7.5 (Ocneanu rigidity for associators in ribbon categories).
Let $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ be a ribbon category over $\mathbb{C}$. Then any deformation $\widetilde{c}$ of $c$ is trivial.
We now reproduce the standard algebro-geometric argument in [12, 9.1.4] to deduce triviality of continuous deformations of associators in ribbon categories. Let us fix a $\mathbb{C}$-linear category $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,\beta,\theta)$ where $\otimes:\mathrm{Vect}_{\Lambda}\times\mathrm{Vect}_{\Lambda}\rightarrow\mathrm{Vect}_{\Lambda}$ is bilinear, $\beta:-_{1}\otimes-_{2}\rightarrow-_{2}\otimes-_{1}$ and $\theta:\mathrm{id}\rightarrow\mathrm{id}$ are linear. Then the data of an associator $c$ for $\otimes$ which makes $\mathcal{C}$ into a ribbon category is a family of linear maps satisfying algebraic conditions. Hence such associators are parametrized by an affine $\mathbb{C}$-scheme of finite type, that we denote $\mathrm{Ass}(\mathcal{C})$.
@@
Let $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ be a ribbon category over $\mathbb{C}$. Then for any continuous family of associators $(c_{t})_{t\in(-\epsilon,\epsilon)}$ with $c_{t}=c_{0}$, there exists a continuous family $(f_{t})_{t\in(-\epsilon,\epsilon)}$ of elements in $\mathrm{Aut}_{\beta}(\otimes)$ such that $f_{0}=\mathrm{id}$ and for all $t$, $c_{t}=f_{t}\cdot c$.
@@
As all the $\mathrm{Aut}_{\beta}(\otimes)$-orbits are open, they are also all closed. Hence the family $(c_{t})_{t\in(-\epsilon,\epsilon)}$ lies in the orbit of $c$. Let $c\in\mathrm{Ass}(\mathcal{C})(\mathbb{C})$, $O_{c}$ its orbit and $G_{c}\subset\mathrm{Aut}_{\beta}(\otimes)$ the stabilizer. Then the associated reduced subscheme $O_{c}^{\mathrm{red}}$ is isomorphic to the quotient $\mathrm{Aut}_{\beta}(\otimes)/G_{c}$. Analytically, $\mathrm{Aut}_{\beta}(\otimes)/G_{c}$ is the quotient of a Lie group by a subgroup. So the action $\mathrm{Aut}_{\beta}(\otimes)\rightarrow O_{c}^{\mathrm{red}}=\mathrm{Aut}_{\beta}(\otimes)/G_{c}$ is a Serre fibration and we can lift the map $c_{t}:(-\epsilon,\epsilon)\rightarrow O_{c}^{\mathrm{red}}$ to $\mathrm{Aut}_{\beta}(\otimes)$. ∎
Let us now turn to infinitesimal automorphisms of ribbon categories.
@@
Let $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ be a ribbon category. An infinitesimal automorphism of $\mathcal{C}$ is a functor of ribbon category $(\mathrm{id}_{\Lambda},\mathrm{id}+\epsilon k):\mathcal{C}_{\mathbb{C}[\epsilon]}\rightarrow\mathcal{C}_{\mathbb{C}[\epsilon]}$. Similarly for fusion categories.
@@
As any trivial infinitesimal fusion automorphism is also a ribbon automorphism, we deduce the following.
@@
Any infinitesimal automorphism of a ribbon category over $\mathbb{C}$ is trivial.
As an application, we prove that all weight functions on a modular functor $\mathcal{V}$ are exact when $\mathrm{C}_{\mathcal{V}}$ is ribbon.
@@
Let $w$ be a weight function on $\mathcal{V}$. Then we can define an infinitesimal automorphism $(\mathrm{id}_{\Lambda},\mathrm{id}+\epsilon k):\mathrm{C}_{\mathcal{V},\mathbb{C}[\epsilon]}\rightarrow\mathrm{C}_{\mathcal{V},\mathbb{C}[\epsilon]}$ by asking that for all $\lambda,\mu,\nu\in\Lambda$, $k$ is on $\hom([\lambda],[\mu]\otimes[\nu])$ multiplication by $w_{0}(\lambda;\mu,\nu)$. Now as $\mathrm{C}_{\mathcal{V}}$ is ribbon (by assumption), Ocneanu rigidity (7.10) implies that $(\mathrm{id}_{\Lambda},\mathrm{id}+\epsilon k)$ is trivial and thus that there exists $f:\Lambda\rightarrow\mathbb{Q}$ with $f(0)=0$ such that for all $\lambda,\mu,\nu\in\Lambda$, $w_{0}(\lambda;\mu,\nu)=f(\mu)+f(\nu)-f(\lambda)$. By the gluing axiom, for any $g$ and $\underline{\lambda}$, $w_{g}(\underline{\lambda})=\frac{1}{2}\sum_{i}f(\lambda_{i})$. ∎
Now, by the same algebro-geometric argument as for deformations of associators (7.7 and above), we deduce from Corollary 7.10 that any continuous family of deformations of the identity ribbon automorphism can be trivialized.
@@
Let $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ be a ribbon category, and $(f_{t})_{t\in(-\epsilon,\epsilon)}$ a continuous family of automorphisms of $\mathcal{C}$ with $f_{0}=\mathrm{id}$. Then there exists a continuous family of maps $(\delta_{t}:\Lambda\rightarrow\mathbb{C}^{*})_{t\in(-\epsilon,\epsilon)}$ with $\delta(0,\cdot)\equiv 1$ such that for any $t$ and $\lambda,\mu,\nu\in\Lambda$, $f_{t}$ is on $\hom([\lambda],[\mu]\otimes[\nu])$ multiplication by $\delta_{t}(\mu)\delta_{t}(\nu)\delta_{t}(\lambda)^{-1}$.

### v2

@@ 5.5. Decomposition complex, following Bakalov-Kirillov
Let $\mathcal{V}$ be a modular functor with associated ribbon fusion category $\mathrm{C}_{\mathcal{V}}$. Then for colors $\lambda_{1},\dotsc,\lambda_{n}\in\Lambda$, we have natural131313with respect to $\mathcal{V}$. isomorphisms:
@@
Let $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ be $2$ modular functors in $\mathfrak{mod}_{c}$ with respective ribbon fusion categories $\mathrm{C}_{\mathcal{V}_{1}}$ and $\mathrm{C}_{\mathcal{V}_{2}}$. Let $G:\mathrm{C}_{\mathcal{V}_{1}}\rightarrow\mathrm{C}_{\mathcal{V}_{2}}$ be a morphism. For a colored surface $\Sigma=(S,\underline{\lambda})$, and a decomposition $D=(\Gamma,(\phi_{S_{i}})_{i})$ of $S$, we define an isomorphism of $R$-modules $G_{D}:\mathcal{V}_{1}(\Sigma)\rightarrow\mathcal{V}_{2}(\Sigma)$ as follows. For $k=1,2$, the gluing axiom along $\Gamma$ expresses $\mathcal{V}_{1}(\Sigma)$ as a direct sum of tensor powers of terms of the form $\mathcal{V}_{k}(S_{i},\underline{\mu})$ or $\mathcal{V}_{k}(S_{0}^{2},\nu,\nu^{\dagger})$. Now we have isomorphisms:
@@
To prove fullness of $\mathfrak{mod}_{c}\rightarrow\mathfrak{Rib}$, we will need that $G_{D}$ is independent of the decomposition $D$ (see Section 5.6). For this, by Theorem 5.16, it is sufficient to prove that $G_{D}$ is invariant under $Z$, $B$, $F$ and $S$-moves. So we need to show that if $D$ and $D^{\prime}$ differ by a move, then the associated decompositions of $\mathcal{V}(\Sigma)$ as in Definition 5.17 are related by isomorphisms depending only on $\mathrm{C}_{\mathcal{V}}$ and natural in $\mathcal{V}$. Below, we explain the invariance under $Z$, $B$, $F$-moves. Invariance under $S$-moves will be the subject of the next subsection 5.6.
@@ 5.6. Higher genus and curve operators, following Andersen-Ueno
Andersen and Ueno prove this for $R=\mathbb{C}$. However their proof works without changes for $R$ any field of characteristic 00.
It consists in showing that a matrix $C$ depending on the data of the modular functor is invertible [2, 5.2]. Assume by contradiction that the statement is false for some $\mathcal{V}$ defined over some ring $R$ all of whose quotients have characteristic 00. Then $\det(C)$ is included in a maximal ideal $\mathfrak{m}$ of $R$. Hence the reduction of $\mathcal{V}$ over $R/\mathfrak{m}$ has non-invertible $C$-matrix. However $R/\mathfrak{m}$ is a field, so this contradicts Andersen and Ueno’s result. ∎
Remark 5.20.
Another proof for $R$ a field of characteristic 00 is provided by the fact that the ribbon category associated to a modular functor is modular (see discussion in the proof of Theorem 5.11). Indeed, the matrix $C$ in Andersen and Ueno’s paper is invertible when the $S$-matrix is.
We can now conclude on the fullness of $\mathfrak{mod}_{c}\rightarrow\mathfrak{Rib}$.
@@
Let $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ be two modular functors with respective ribbon fusion categories $\mathrm{C}_{\mathcal{V}_{1}}$ and $\mathrm{C}_{\mathcal{V}_{2}}$. Choose a morphisms $G:\mathrm{C}_{\mathcal{V}_{1}}\rightarrow\mathrm{C}_{\mathcal{V}_{2}}$. Then for each decompositions $D$ of a surface $\Sigma$, Definition 5.17 provides an isomorphism $G_{D}:\mathcal{V}_{1}(\Sigma)\rightarrow\mathcal{V}_{2}(\Sigma)$. If we show that $G_{D}$ is independent of the decompositions $D$, then setting, for each colored surface $\Sigma$, $G_{\Sigma}:=G_{D}$, the family $(G_{\Sigma})_{\Sigma}$ provides an isomorphism of modular functor $\widetilde{G}:\mathcal{V}_{1}\rightarrow\mathcal{V}_{2}$ lifting $G:\mathrm{C}_{\mathcal{V}_{1}}\rightarrow\mathrm{C}_{\mathcal{V}_{2}}$.
@@ 6.1. Higgs bundles and local systems
Let $X$ be a smooth Deligne-Mumford stack. Then we have equivalences between the $\mathbb{C}$-linear symmetric tensor categories 141414i.e. with a tensor product and duals. of:
@@ A Higgs bundle $(\mathcal{E},\theta)$ on a proper smooth DM stack $X$ is of harmonic type if:
Let $\Lambda$ be a set of colors and $r,s\geq 1$ integers. Then the definition of a modular functor of Higgs bundles with level $(r,s)$ is obtained from that of a (non-anomalous) complex modular functor (2.30, 2.45) by replacing the $\mathbb{C}$-local systems $\mathcal{V}_{g}(\underline{\lambda})$ by Higgs bundles $(\mathcal{V}_{g}(\underline{\lambda}),\theta_{g}(\underline{\lambda}))$.
Similarly for genus 00 modular functors, ribbon and braided functors.
From the correspondence, its functoriality and Theorem 3.8 we immediately get the following.
@@
complex modular functors;
@@
Similarly for genus 00 modular functors, ribbon functors and braided functors.
@@ 6.2. Systems of Hodge bundles
Hence we can split $\nabla$ into $\overline{\theta}+\partial_{E}+\overline{\partial}_{E}+\theta$ according to the decomposition above. Then $(E,\overline{\partial}_{E},\theta)$ is the associated system of Hodge bundles, and any polarization for $(E,\nabla)$ is also one for $(E,\overline{\partial}_{E},\theta)$. To see this, notice that for any polarization $h$, the metric $K=\sum_{p+q=m}\mathrm{sign}(h_{\mid E^{p,q}})h_{\mid E^{p,q}}$ is pluri-harmonic.
Let $(\mathcal{E},\theta)$ be a system of Hodge bundles on $(\mathcal{E},\theta)$. Denote by $E$ the underlying $C^{\infty}$ bundle and by ¯$\overline{\partial}_{E}$ the Dolbeault operator. Choose a polarization $h$. Now let $\partial_{E}$ be the unique $(1,0)$-connection such that $\partial_{E}+\overline{\partial}_{E}$ preserves $h$, and let $\overline{\theta}$ be the adjoint of $\theta$ for $h$. Then $(E,\nabla=\overline{\theta}+\partial_{E}+\overline{\partial}_{E}+\theta)$ is the associated CVHS (with real bidegrees) and $h$ polarizes it. Moreover $\nabla$ is independent of the choice of the polarization $h$. Again, to see this, notice that the metric $K=\sum_{p+q=m}\mathrm{sign}(h_{\mid E^{p,q}})h_{\mid E^{p,q}}$ is Hermitian-Yang-Mills.
@@
Let $\mathcal{V}$ be a complex modular functor. Let $\Lambda$ be its set of colors and $(\mathcal{V},\theta)$ the associated modular functor of Higgs bundles (6.15). Choose a weight function $w$ on $\mathcal{V}$ and some $t\in\mathbb{R}_{+}^{*}\setminus\{1\}$. Then the maps $i$ and $r$ of Lemma 6.20 induce maps:
@@ 7. Ocneanu rigidity
7.1. Deformations of associators and endofunctors of ribbon fusion categories
@@
Note that our definition of morphisms of fusion and ribbon fusion categories is more restrictive than the usual one.
@@
If $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ is a ribbon fusion category over $\mathbb{C}$, a deformation of $c$ is an associator $\widetilde{c}$ such that $(\mathrm{Vect}_{\Lambda}^{{\mathbb{C}[\epsilon]}},\otimes_{{\mathbb{C}[\epsilon]}},\widetilde{c},\beta_{\mathbb{C}[\epsilon]},\theta_{\mathbb{C}[\epsilon]})$ is a braided category with a twist. Such a deformation is trivial if there exists a pair $(\mathrm{id}_{\Lambda},(\mathrm{id}+\epsilon\phi))$ as above that is an isomorphism compatible with the braiding and twist. Similarly for braided fusion categories over $\mathbb{C}$.
Note that any deformation $\widetilde{c}$ of $c$ is always of the form $c(\mathrm{id}+\epsilon d)$ for $d:(-\otimes-)\otimes-\rightarrow(-\otimes-)\otimes-$ a natural transformation.
@@
We now briefly explain how this result for fusion categories can be extended to ribbon fusion categories. A quick proof of this is explained in [13, Rmk. 2.33]. Here we give another proof, and take this as an opportunity to better understand the meaning of Theorem 7.4. Fix a ribbon fusion category $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$. Let $C^{3}(\mathcal{C})$ be the vector space of all natural transformations $d:(-\otimes-)\otimes-\rightarrow(-\otimes-)\otimes-$, let $Z^{3}(\mathcal{C})$ be the subspace of the $d$ such that $c(\mathrm{id}+d)$ is an associator for $(\mathrm{Vect}_{\Lambda}^{{\mathbb{C}[\epsilon]}},\otimes_{{\mathbb{C}[\epsilon]}})$, and let $Z^{3}(\mathcal{C},\beta)$ be the subspace of the $d$ such that $c(\mathrm{id}+d)$ is an associator compatible with the braiding $\beta_{\mathbb{C}[\epsilon]}$.
@@
Ocneanu rigidity (7.4) is equivalent to the statement that $\partial$ is surjective. The map $\partial$ is a differential in the proper Davydov-Yetter complex of $\mathrm{id}_{\mathcal{C}}$. This complex is quasi-isomorphic to the usual Davydov-Yetter complex (see [16, (47) and B.3]). Ocneanu rigidity is usually stated as the acyclicity of this latter complex (see for example [13, 7.22, 9.1.3]).
@@
Corollary 7.5 (Ocneanu rigidity for associators in ribbon and braided fusion categories ([12, 2.33])).
Let $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ be a ribbon fusion category over $\mathbb{C}$. Then any deformation $\widetilde{c}$ of $c$ is trivial. Similarly for braided fusion categories over $\mathbb{C}$.
We now reproduce the standard algebro-geometric argument in [12, 9.1.4] to deduce triviality of continuous deformations of associators in ribbon fusion categories. Let us fix a $\mathbb{C}$-linear category $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,\beta,\theta)$ where $\otimes:\mathrm{Vect}_{\Lambda}\times\mathrm{Vect}_{\Lambda}\rightarrow\mathrm{Vect}_{\Lambda}$ is bilinear, $\beta:-_{1}\otimes-_{2}\rightarrow-_{2}\otimes-_{1}$ and $\theta:\mathrm{id}\rightarrow\mathrm{id}$ are linear. Then the data of an associator $c$ for $\otimes$ which makes $\mathcal{C}$ into a ribbon category is a family of linear maps satisfying algebraic conditions. Hence such associators are parametrized by an affine $\mathbb{C}$-scheme of finite type, that we denote $\mathrm{Ass}(\mathcal{C})$.
@@
Let $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ be a ribbon fusion category over $\mathbb{C}$. Then for any continuous family of associators $(c_{t})_{t\in(-1,1)}$ with $c_{0}=c$, there exists a continuous family $(f_{t})_{t\in(-1,1)}$ of elements in $\mathrm{Aut}_{\beta}(\otimes)$ such that $f_{0}=\mathrm{id}$ and for all $t$, $c_{t}=f_{t}\cdot c$. The same result holds in the braided fusion case.
@@
As all the $\mathrm{Aut}_{\beta}(\otimes)$-orbits are open, they are also all closed. Hence the family $(c_{t})_{t\in(-1,1)}$ lies in the orbit of $c$. Let $c\in\mathrm{Ass}(\mathcal{C})(\mathbb{C})$, $O_{c}$ its orbit and $G_{c}\subset\mathrm{Aut}_{\beta}(\otimes)$ the stabilizer. Then the associated reduced subscheme $O_{c}^{\mathrm{red}}$ is isomorphic to the quotient $\mathrm{Aut}_{\beta}(\otimes)/G_{c}$. Analytically, $\mathrm{Aut}_{\beta}(\otimes)/G_{c}$ is the quotient of a Lie group by a subgroup. So the action $\mathrm{Aut}_{\beta}(\otimes)\rightarrow O_{c}^{\mathrm{red}}=\mathrm{Aut}_{\beta}(\otimes)/G_{c}$ is a Serre fibration and we can lift the map $c_{t}:(-1,1)\rightarrow O_{c}^{\mathrm{red}}$ to $\mathrm{Aut}_{\beta}(\otimes)$. ∎
Let us now turn to infinitesimal automorphisms of ribbon fusion categories.
@@
Let $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ be a ribbon fusion category. An infinitesimal automorphism of $\mathcal{C}$ is a functor of ribbon categories $(\mathrm{id}_{\Lambda},\mathrm{id}+\epsilon k):\mathcal{C}_{\mathbb{C}[\epsilon]}\rightarrow\mathcal{C}_{\mathbb{C}[\epsilon]}$. Similarly for fusion categories and braided fusion categories.
@@
As any trivial infinitesimal fusion automorphism is also a braided/ribbon automorphism, we deduce the following.
@@
Any infinitesimal automorphism of a braided (or ribbon) fusion category over $\mathbb{C}$ is trivial.
As an application, we prove that all weight functions on a modular functor are exact.
@@
Let $w$ be a weight function on $\mathcal{V}$. Then we can define an infinitesimal automorphism $(\mathrm{id}_{\Lambda},\mathrm{id}+\epsilon k):\mathrm{C}_{\mathcal{V},\mathbb{C}[\epsilon]}\rightarrow\mathrm{C}_{\mathcal{V},\mathbb{C}[\epsilon]}$ by asking that for all $\lambda,\mu,\nu\in\Lambda$, $k$ is on $\hom([\lambda],[\mu]\otimes[\nu])$ multiplication by $w_{0}(\lambda;\mu,\nu)$. Now, Ocneanu rigidity (7.10) implies that $(\mathrm{id}_{\Lambda},\mathrm{id}+\epsilon k)$ is trivial and thus that there exists $f:\Lambda\rightarrow\mathbb{Q}$ with $f(0)=0$ such that for all $\lambda,\mu,\nu\in\Lambda$, $w_{0}(\lambda;\mu,\nu)=f(\mu)+f(\nu)-f(\lambda)$. By the gluing axiom, for any $g$ and $\underline{\lambda}$, $w_{g}(\underline{\lambda})=\frac{1}{2}\sum_{i}f(\lambda_{i})$. ∎
Now, by the same algebro-geometric argument as for deformations of associators (7.7 and above), we deduce from Corollary 7.10 that any continuous family of deformations of the identity braided/ribbon automorphism can be trivialized.
@@
Let $\mathcal{C}=(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ be a ribbon fusion category, and $(f_{t})_{t\in(-\epsilon,\epsilon)}$ a continuous family of automorphisms of $\mathcal{C}$ with $f_{0}=\mathrm{id}$. Then there exists a continuous family of maps $(\delta_{t}:\Lambda\rightarrow\mathbb{C}^{*})_{t\in(-\epsilon,\epsilon)}$ with $\delta(0,\cdot)\equiv 1$ such that for any $t$ and $\lambda,\mu,\nu\in\Lambda$, $f_{t}$ is on $\hom([\lambda],[\mu]\otimes[\nu])$ multiplication by $\delta_{t}(\mu)\delta_{t}(\nu)\delta_{t}(\lambda)^{-1}$. The same holds for braided fusion categories.